# ScoreAdv: Score-based Targeted Generation of Natural Adversarial Examples via Diffusion Models

### v1

@@ Abstract
Despite the success of deep learning across various domains, it remains vulnerable to adversarial attacks. Although many existing adversarial attack methods achieve high success rates, they typically rely on $\ell_{p}$-norm perturbation constraints, which do not align with human perceptual capabilities. Consequently, researchers have shifted their focus toward generating natural, unrestricted adversarial examples (UAEs). GAN-based approaches suffer from inherent limitations, such as poor image quality due to instability and mode collapse. Meanwhile, diffusion models have been employed for UAE generation, but they still rely on iterative PGD perturbation injection, without fully leveraging their central denoising capabilities. In this paper, we introduce a novel approach for generating UAEs based on diffusion models, named ScoreAdv. This method incorporates an interpretable adversarial guidance mechanism to gradually shift the sampling distribution towards the adversarial distribution, while using an interpretable saliency map to inject the visual information of a reference image into the generated samples. Notably, our method is capable of generating an unlimited number of natural adversarial examples and can attack not only classification models but also retrieval models. We conduct extensive experiments on ImageNet and CelebA datasets, validating the performance of ScoreAdv across ten target models in both black-box and white-box settings. Our results demonstrate that ScoreAdv achieves state-of-the-art attack success rates and image quality. Furthermore, the dynamic balance between denoising and adversarial perturbation enables ScoreAdv to remain robust even under defensive measures.
Codehttps://github.com/6tonystark6/ScoreAdv
Introduction
@@
Recent years have witnessed the remarkable application of deep learning in various domains, including image classification (Cheng, Yap, and Wen 2023), autonomous driving (Feng et al. 2023), medical image analysis (Zhang, Xie, and Chen 2023), etc. However, studies such as (Szegedy et al. 2014) have demonstrated that deep learning models are highly susceptible to adversarial examples—subtle, imperceptible perturbations that mislead models into wrong predictions with high confidence (Huang and Shen 2025). The transferability of adversarial examples poses an even greater security risk, as adversarial samples crafted for one target model can be effectively transferred to different models, potentially compromising the security of multiple systems.
@@
Many previous adversarial attacks have been based on $\ell_{p}$-norm constraints (Dong et al. 2019; Zhao et al. 2024) to bound perturbation strength. However, recent research (Zhao, Liu, and Larson 2020) has shown that $\ell_{p}$-norm poorly aligns with human perceptual similarity. As shown in Figure 1, attacks based on $\ell_{p}$-norm, despite having a small $\ell_{\infty}$ value, are more easily noticeable to the human eye.
@@
In contrast to attacks based on $\ell_{p}$-norm perturbations, this paper investigates unrestricted adversarial example generation, where the generated samples are not constrained by any norm limitations, but instead optimize the input noise vector directly. Such perturbations typically exhibit large-scale, semantically coherent patterns that enhance attack success rates while remaining imperceptible (Jia et al. 2022).
The concept of UAE was first introduced by Song et al. (Song et al. 2018), who employed a generative adversarial network (GAN) with an auxiliary classifier to directly generate adversarial samples. However, GAN-based methods suffer from instability, mode collapse, and limited interpretability, often resulting in lower-quality outputs. In contrast, diffusion models (Ho, Jain, and Abbeel 2020) offer more stable training, better image quality, and broader distribution coverage (Dhariwal and Nichol 2021), making them more suitable for generating natural adversarial images.
The iterative denoising process in diffusion models naturally serves as a robust mechanism, capable of mitigating adversarial noise while retaining attack efficacy. (Dhariwal and Nichol 2021) also highlighted that conditions can be incorporated into the denoising process to guide the generation. This characteristic enables us to carefully introduce adversarial information during the denoising process, thus achieving the generation of natural adversarial samples. However, existing diffusion-based attack methods have not fully exploited this capability. They often rely on PGD-based perturbations at each step (Chen et al. 2023a, b), or lack reference-image conditioning (Dai, Liang, and Xiao 2025), limiting image quality and the full use of diffusion’s potential.
In this paper, we propose a novel and interpretable UAE generation method based on diffusion models, called ScoreAdv. Compared to previous diffusion-based adversarial attack methods, ScoreAdv 1) introduces an interpretable adversarial perturbation method at each diffusion step, gradually shifting the sampling distribution towards the adversarial distribution. 2) Proposes using ScoreCAM to highlight the salient regions of a reference image, enabling the generation of adversarial examples that are consistent with the reference image. 3) Iteratively optimizes the initial sampling noise to incorporate both the target semantic information and the visual features of the reference image, improving the adversarial sample’s alignment with the intended objectives.
@@
    We propose ScoreAdv, a training-free, unrestricted adversarial example generation framework based on diffusion models. ScoreAdv can generate an unlimited number of natural adversarial examples, targeting not only classification but also recognition models.
@@
    Comprehensive experiments on ten target models demonstrate that ScoreAdv outperforms existing methods in both black-box and white-box settings, achieving superior attack success and visual quality. The dynamic interplay between denoising and adversarial perturbation also improves robustness against defenses.
Related Work
@@
Adversarial Attacks
Adversarial attacks are commonly classified as white-box or black-box, depending on the attacker’s knowledge of the target model. White-box attacks assume full access to model architecture, gradients, and parameters, with representative methods including FGSM (Goodfellow, Shlens, and Szegedy 2015), PGD (Madry et al. 2018a), and DeepFool (Moosavi-Dezfooli, Fawzi, and Frossard 2016). In contrast, black-box attacks, which only rely on input-output queries, are more challenging and better reflect real-world conditions. Notable examples include ZOO (Chen et al. 2017) and DeepSearch (Zhang, Chowdhury, and Christakis 2020). Attacks can also be untargeted, aiming to misclassify into any incorrect label, or targeted, forcing misclassification into a specific class, often posing greater threats. Many black-box attacks leverage transferability (Papernot, McDaniel, and Goodfellow 2016; Dong et al. 2018), applying white-box techniques on surrogate models to generate transferable adversarial examples.
Unrestricted Adversarial Attacks
Due to the inadequacy of $\ell_{p}$-norm in capturing perceptual similarity in RGB space, researchers have increasingly explored unrestricted adversarial attacks. Early works, such as (Zhao, Dua, and Singh 2018), modified GAN latent codes via auxiliary classifiers, but suffered from limited image quality. (Bhattad et al. 2019) perturbed color and texture, while (Yuan et al. 2022) used a color distribution library and a neighborhood search to improve realism. (Jia et al. 2022) generated attribute-based adversarial noise targeting facial features. More recent methods leverage diffusion models for unrestricted attacks. (Dai, Liang, and Xiao 2025) introduced adversarial guidance for stable generation; (Chen et al. 2023a) proposed AdvDiffuser, injecting PGD perturbations during sampling; and (Chen et al. 2025) employed latent-space mapping with content-preserving structures and attention-based distractions to craft adversarial examples.
Diffusion Models
Diffusion models are based on stochastic processes and Markov chains, with a forward process of gradually adding noise and a reverse denoising process. Given an image $\boldsymbol{x}_{0}$, the forward process adds noise over $T$ steps using a schedule $\beta_{1:T}$, generating a increasing noisy image sequence.
@@
Subsequent advances include Improved DDPM (Nichol and Dhariwal 2021), which learns a variance schedule to improve sampling quality and efficiency. (Dhariwal and Nichol 2021) introduced conditional diffusion, incorporating the gradient of cross-entropy loss from a target model into the sampling mean for class-conditioned generation.
@@
Furthermore, classifier-free guidance (Ho and Salimans 2021) enables the incorporation of class information without the need for an additional trained classifier.
@@
Methodology
Formulation
Instead of adding perturbations to clean samples, we aim to leverage a generative model to produce UAEs (Song et al. 2018), yielding more natural adversarial inputs. Given a generative model $\mathcal{G}$, a target model $f$, ground truth label $y$, target label $y_{tar}$, and an optional reference image $\boldsymbol{x}^{ref}$, the generated UAE is defined as:
@@
Algorithm 1 ScoreAdv Diffusion Sampling
0:  Ground truth label $y$, (optional) reference image $\boldsymbol{x}^{ref}$, target label $y_{tar}$ or target image $\boldsymbol{x}_{tar}$, target model $f$, pretrained diffusion model $\epsilon_{\theta}$, reverse generation process timestep $T$, noise optimization times $N$, adversarial guidance scale $s_{a}$, classifier guidance scale $s_{c}$, noise optimization guidance scale $s_{n}$, noise schedule $\beta_{1:T}$;
0:  Adversarial image $\boldsymbol{x}_{adv}$;
1:  $\boldsymbol{x}_{T}\sim\mathcal{N}\left(0,\boldsymbol{I}\right)$, $\boldsymbol{x}_{adv}=\{\}$;
2:  for all $i=1,\ldots,N$ do
3:     for all $t=T,\ldots,1$ do
4:        $\tilde{\boldsymbol{\epsilon}}_{t}=\left(1+s_{c}\right)\epsilon_{\theta}\left(\boldsymbol{x}_{t},y\right)-s_{c}\epsilon_{\theta}(\boldsymbol{x}_{t})$;
5:        $\boldsymbol{\bar{x}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\boldsymbol{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\tilde{\epsilon}_{t}\right)+\sigma_{t}z_{t}$;
6:        if $\boldsymbol{x}^{ref}$ then
7:           $\boldsymbol{m}=\operatorname{ScoreCAM}\left(\boldsymbol{x}^{ref},f,y\right)$;
8:           $\boldsymbol{x}_{t-1}^{ref}\sim\mathcal{N}\left(\sqrt{\overline{\alpha}_{t-1}}\boldsymbol{x}^{ref},\left(1-\overline{\alpha}_{t-1}\right)\boldsymbol{I}\right)$;
9:           $\boldsymbol{x}_{t-1}=\boldsymbol{\tilde{x}}_{t-1}\odot\left(\boldsymbol{1}-\boldsymbol{m}\right)+\boldsymbol{x}_{t-1}^{ref}\odot\boldsymbol{m}$;
10:        end if
11:        $\hat{\boldsymbol{x}}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\boldsymbol{\bar{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t}}\cdot\tilde{\boldsymbol{\epsilon}}_{t-1}\right)$;
12:        $\boldsymbol{g}=\nabla_{\boldsymbol{x}_{t-1}}\log p_{f}\left(y_{tar}|\boldsymbol{\bar{x}}_{t-1}\right)$; if $y_{tar}$
13:        $\boldsymbol{g}=\nabla_{\boldsymbol{\hat{\boldsymbol{x}}_{0}}}\log p_{f}\left(\boldsymbol{x}_{tar}|\hat{\boldsymbol{x}}_{0}\right)$; if not $y_{tar}$
14:        $\boldsymbol{\tilde{x}}_{t-1}=\boldsymbol{\bar{x}}_{t-1}+\sigma_{t}^{2}s_{a}\boldsymbol{g}$;
15:     end for
16:     $\boldsymbol{g}_{n}=\nabla_{\boldsymbol{x}_{0}}\log p_{f}\left(y_{tar}|\boldsymbol{x}_{0}\right)$; if $\boldsymbol{x}^{ref}$
17:     $\boldsymbol{g}_{n}=\nabla_{\boldsymbol{x}_{0}}\log p_{f}\left(\boldsymbol{x}_{tar}|\boldsymbol{x}_{0}\right)$; if not $\boldsymbol{x}^{ref}$
18:     $\boldsymbol{x}_{T}\leftarrow\boldsymbol{x}_{T}+\bar{\sigma}_{T}^{2}s_{n}\boldsymbol{g}_{n}$;
19:     if $f(\boldsymbol{x}_{0})=y_{tar}$ then
20:        $\boldsymbol{x}_{adv}\leftarrow\boldsymbol{x}_{0}$;
21:     end if
22:  end for
Adversarial Perturbation
When using a pre-trained diffusion model, the denoising process gradually transforms noise into a natural image. Inspired by (Dhariwal and Nichol 2021), we found that during the denoising process, the gradient $\nabla_{\boldsymbol{x}_{t}}\log p_{\phi}(y|\boldsymbol{x}_{t})$ can be used to guide the shift of the sampling distribution. Our objective is to generate an image $\boldsymbol{x}_{0}$ with a pre-trained diffusion model $\epsilon_{\theta}(\boldsymbol{x}_{t},y)$, which appears consistent with the target label $y$, but is misclassified by the target model as $y_{tar}\neq y$. We achieve this through adversarial perturbations.
@@
The derivation of Equation (8) is provided in Supplementary Material. Intuitively, a small perturbation is introduced to shift $\bar{\boldsymbol{x}}_{t-1}$ toward the target class $y_{tar}$, increasing the likelihood that the generated sample $\tilde{\boldsymbol{x}}_{t-1}$ is classified as $y_{tar}$.

### v2

@@ Abstract
Despite the success of deep learning across various domains, it remains vulnerable to adversarial attacks. Although many existing adversarial attack methods achieve high success rates, they typically rely on $\ell_{p}$-norm perturbation constraints, which do not align with human perceptual capabilities. Consequently, researchers have shifted their focus toward generating natural, unrestricted adversarial examples (UAEs). GAN-based approaches suffer from inherent limitations, such as poor image quality due to instability and mode collapse. Meanwhile, diffusion models have been employed for UAE generation, but they still rely on iterative PGD perturbation injection, without fully leveraging their central denoising capabilities. In this paper, we introduce a novel approach for generating UAEs based on diffusion models, named ScoreAdv. This method incorporates an interpretable adversarial guidance mechanism to gradually shift the sampling distribution towards the adversarial distribution, while using an interpretable saliency map to inject the visual information of a reference image into the generated samples. Notably, our method is capable of generating an unlimited number of natural adversarial examples and can attack not only classification models but also retrieval models. We conduct extensive experiments on ImageNet and CelebA datasets, validating the performance of ScoreAdv across ten target models in both black-box and white-box settings. Our results demonstrate that ScoreAdv achieves state-of-the-art attack success rates and image quality, while maintaining inference efficiency. Furthermore, the dynamic balance between denoising and adversarial perturbation enables ScoreAdv to remain robust even under defensive measures.
1 Introduction
@@
Recent years have witnessed the remarkable application of deep learning in various domains, including image classification [8], autonomous driving [15], medical image analysis [61], face recognition [40], etc. However, studies such as those of [45] have demonstrated that deep learning models are highly susceptible to adversarial examples, which can mislead them into making incorrect decisions, thus compromising their security in real-world applications. Adversarial examples refer to subtle, imperceptible perturbations added to clean examples that cause the target model to make erroneous predictions with high confidence while remaining undetectable to the human eye. The transferability of adversarial examples poses an even greater security risk, as adversarial samples crafted for one target model can be effectively transferred to different models, potentially compromising the security of multiple systems.
@@
Many previous adversarial attacks have been based on $\ell_{p}$-norm constraints [14, 62] to bound perturbation strength. However, recent research [64] has shown that $\ell_{p}$-norm poorly aligns with human perceptual similarity. As shown in Figure 1, attacks based on $\ell_{p}$-norm, despite having a small $\ell_{\infty}$ value, are more easily noticeable to the human eye.
@@
In contrast to attacks based on $\ell_{p}$-norm perturbations, this paper investigates unrestricted adversarial example generation, where the generated samples are not constrained by any norm limitations, but instead optimize the input noise vector directly. Such perturbations typically exhibit large-scale, semantically coherent patterns that enhance attack success rates while remaining imperceptible [22].
The concept of UAE was first introduced by [44], who employed a generative adversarial network (GAN) with an auxiliary classifier to directly generate adversarial samples. However, GAN-based methods suffer from instability, mode collapse, and limited interpretability, often resulting in lower-quality outputs. In contrast, diffusion models [20] offer more stable training, better image quality, and broader distribution coverage [12], making them more suitable for generating natural adversarial images.
The iterative denoising process in diffusion models naturally serves as a robust mechanism, capable of mitigating adversarial noise while retaining attack efficacy. [12] also highlighted that conditions can be incorporated into the denoising process to guide the generation. This characteristic enables us to carefully introduce adversarial information during the denoising process, thus achieving the generation of natural adversarial samples. However, existing diffusion-based attack methods have not fully exploited this capability. They often rely on PGD-based perturbations at each step [6, 7], or lack reference-image conditioning [11], limiting image quality and the full use of diffusion’s potential.
Generating natural-looking adversarial examples from diffusion priors poses two linked challenges: (1) steering the stochastic sampling trajectory toward adversarial regions of the data manifold without destroying realism, and (2) doing so in a way that generalizes across classifiers and recognition systems under realistic constraints.
In this paper, we propose ScoreAdv, a novel and interpretable UAE generation method based on diffusion models, which addresses both challenges by moving the adversarial mechanism inside the diffusion posterior instead of treating perturbation as an external add-on. Compared to previous diffusion-based adversarial attack methods, ScoreAdv (I) internalize stepwise, score-based adversarial guidance that reweights the denoising posterior at every sampling step, (II) fuse semantic structure via ScoreCAM-derived masks computed on a surrogate model to preserve fine-grained content, and (III) perform lightweight noise optimization to align high-variance latents with adversarial objectives without retraining the model. This unique combination yields a training-free, model-agnostic pipeline that attacks both classifiers and recognition systems, delivers strong black-box transfer, and keeps inference costs similar to the original diffusion model.
@@
    We propose ScoreAdv, a training-free, unrestricted adversarial example generation framework based on diffusion models. It can generate an unlimited number of natural adversarial images, and its internalized score-based guidance yields a unified mechanism for both classification and recognition attacks while preserving realism.
@@
    Comprehensive experiments on ten target models demonstrate that ScoreAdv outperforms existing methods in both black-box and white-box settings, achieving superior attack success and visual quality, while keeping inference highly efficient. The dynamic interplay between denoising and adversarial perturbation also improves robustness against defenses.
2 Related Work
@@
2.1 Adversarial Attacks
Adversarial attacks are commonly classified as white-box or black-box, depending on the attacker’s knowledge of the target model. White-box attacks assume full access to model architecture, gradients, and parameters, with representative methods including FGSM [16], PGD [33], and DeepFool [34]. In contrast, black-box attacks, which only rely on input-output queries, are more challenging and better reflect real-world conditions. Notable examples include ZOO [4] and DeepSearch [59]. Attacks can also be untargeted, aiming to misclassify into any incorrect label, or targeted, forcing misclassification into a specific class, often posing greater threats. Many black-box attacks leverage transferability [39, 13], applying white-box techniques on surrogate models to generate transferable adversarial examples.
2.2 Unrestricted Adversarial Attacks
Due to the inadequacy of $\ell_{p}$-norm in capturing perceptual similarity in RGB space, researchers have increasingly explored unrestricted adversarial attacks. Early works, such as [63], modified GAN latent codes via auxiliary classifiers, but suffered from limited image quality. [1] perturbed color and texture, while [57] used a color distribution library and a neighborhood search to improve realism. [22] generated attribute-based adversarial noise targeting facial features. More recent methods leverage diffusion models for unrestricted attacks. [11] introduced adversarial guidance for stable generation; [6] proposed AdvDiffuser, injecting PGD perturbations during sampling; and [3] employed latent-space mapping with content-preserving structures and attention-based distractions to craft adversarial examples.
2.3 Diffusion Models
Diffusion models are based on stochastic processes and Markov chains, involving a forward process of gradually adding noise and a reverse denoising process. Given an image $\boldsymbol{x}_{0}$, a total of $T$ diffusion steps, and a noise schedule $\beta_{1:T}$, the forward process uses a fixed Markov chain to iteratively add Gaussian noise, producing a sequence of increasingly degraded images.
@@ 2.3 Diffusion Models
Subsequent advances include Improved DDPM [37], which learns a variance schedule to improve sampling quality and efficiency. [12] introduced conditional diffusion, incorporating the gradient of cross-entropy loss from a target model into the sampling mean for class-conditioned generation.
@@
Furthermore, classifier-free guidance [19] enables the incorporation of class information without the need for an additional trained classifier.
@@
3 Methodology
3.1 Formulation
Instead of adding perturbations to clean samples, we aim to leverage a generative model to produce UAEs [44], yielding more natural adversarial inputs. Given a generative model $\mathcal{G}$, a target model $f$, ground truth label $y$, target label $y_{tar}$, and an optional reference image $\boldsymbol{x}^{ref}$, the generated UAE is defined as:
@@ 3.1 Formulation
3.2 Adversarial Perturbation
When using a pre-trained diffusion model, the denoising process gradually transforms noise into a natural image. Inspired by [12], we found that during the denoising process, the gradient $\nabla_{\boldsymbol{x}_{t}}\log p_{\phi}(y|\boldsymbol{x}_{t})$ can be used to guide the shift of the sampling distribution. Our objective is to generate an image $\boldsymbol{x}_{0}$ with a pre-trained diffusion model $\epsilon_{\theta}(\boldsymbol{x}_{t},y)$, which appears consistent with the target label $y$, but is misclassified by the target model as $y_{tar}\neq y$. We achieve this through adversarial perturbations.
@@ 3.2 Adversarial Perturbation
The derivation of Equation (8) is provided in Supplementary Material 6.1. Intuitively, a small perturbation is introduced to shift $\bar{\boldsymbol{x}}_{t-1}$ toward the target class $y_{tar}$, increasing the likelihood that the generated sample $\tilde{\boldsymbol{x}}_{t-1}$ is classified as $y_{tar}$.
Input: Ground truth class label $y$, (optional) reference image $\boldsymbol{x}^{ref}$, target label $y_{tar}$ or target image $\boldsymbol{x}_{tar}$, target model $f$, pretrained diffusion model $\epsilon_{\theta}$, reverse generation process timestep $T$, noise optimization times $N$, adversarial guidance scale $s_{a}$, classifier guidance scale $s_{c}$, noise optimization guidance scale $s_{n}$, noise schedule $\beta_{1:T}$
Output: Adversarial image $\boldsymbol{x}_{adv}$
1
[…]
3 $\boldsymbol{x}_{adv}=\{\}$;
4 for $i=1,\ldots,N$ do
5 for $t=T,\ldots,1$ do
    // classifier-free sampling
6    $\tilde{\boldsymbol{\epsilon}}_{t}=\left(1+s_{c}\right)\epsilon_{\theta}\left(\boldsymbol{x}_{t},y\right)-s_{c}\epsilon_{\theta}(\boldsymbol{x}_{t})$;
7    $\boldsymbol{\bar{x}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\boldsymbol{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\tilde{\epsilon}_{t}\right)+\sigma_{t}z_{t}$;
    // if use $\boldsymbol{x}^{ref}$
8    if $\boldsymbol{x}^{ref}$ then
9       $\boldsymbol{m}=\operatorname{ScoreCAM}\left(\boldsymbol{x}^{ref},f,y\right)$;
10       $\boldsymbol{x}_{t-1}^{ref}\sim\mathcal{N}\left(\sqrt{\overline{\alpha}_{t-1}}\boldsymbol{x}^{ref},\left(1-\overline{\alpha}_{t-1}\right)\boldsymbol{I}\right)$;
11       $\boldsymbol{\bar{x}}_{t-1}=\boldsymbol{\tilde{x}}_{t-1}\odot\left(\boldsymbol{1}-\boldsymbol{m}\right)+\boldsymbol{x}_{t-1}^{ref}\odot\boldsymbol{m}$;
12      end if
    // adversarial gradient
13    $\hat{\boldsymbol{x}}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t-1}}}\left(\boldsymbol{\bar{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}\cdot\tilde{\boldsymbol{\epsilon}}_{t-1}\right)$;
14    if $y_{tar}$ then $\boldsymbol{g}=\nabla_{\boldsymbol{x}_{t-1}}\log p_{f}\left(y_{tar}|\boldsymbol{\bar{x}}_{t-1}\right)$;
15    else $\boldsymbol{g}=\nabla_{\boldsymbol{\hat{\boldsymbol{x}}_{0}}}\log p_{f}\left(\boldsymbol{x}_{tar}|\hat{\boldsymbol{x}}_{0}\right)$;
16    $\boldsymbol{\tilde{x}}_{t-1}=\boldsymbol{\bar{x}}_{t-1}+\sigma_{t}^{2}s_{a}\boldsymbol{g}$;
17
18   end for
 // noise gradient
19 if $y_{tar}$ then $\boldsymbol{g}_{n}=\nabla_{\boldsymbol{x}_{0}}\log p_{f}\left(y_{tar}|\boldsymbol{x}_{0}\right)$;
20 else $\boldsymbol{g}_{n}=\nabla_{\boldsymbol{x}_{0}}\log p_{f}\left(\boldsymbol{x}_{tar}|\boldsymbol{x}_{0}\right)$;
21 $\boldsymbol{x}_{T}\leftarrow\boldsymbol{x}_{T}+\bar{\sigma}_{T}^{2}s_{n}\boldsymbol{g}_{n}$;
 // reserve if attack successful
22 if $f(\boldsymbol{x}_{0})=y_{tar}$ then $\boldsymbol{x}_{adv}\leftarrow\boldsymbol{x}_{0}$;
23
24 end for
Algorithm 1 ScoreAdv Diffusion Sampling
Image Recognition.
While the above adversarial perturbation targets classification tasks, our method also extends to image recognition. Given a target image $\boldsymbol{x}_{tar}$ and a similarity model $f(\cdot,\cdot)$, we apply the same procedure to obtain $\bar{\boldsymbol{x}}_{t-1}$, and then sample $\tilde{\boldsymbol{x}}_{t-1}$ as follows:
$\hat{\boldsymbol{x}}_{0}$ $=\frac{1}{\sqrt{\bar{\alpha}_{t-1}}}\left(\bar{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}\cdot\tilde{\boldsymbol{\epsilon}}_{t-1}\right),$
where $\hat{\boldsymbol{x}}_{0}$ is the predicted clean image from $\bar{\boldsymbol{x}}_{t-1}$, and $\tilde{\boldsymbol{\epsilon}}_{t-1}$ is the predicted noise at step t1t-1. We measure similarity between $\boldsymbol{x}_{tar}$ and $\hat{\boldsymbol{x}}_{0}$ rather than $\bar{\boldsymbol{x}}_{t-1}$, since the latter contains significant noise. As the target model is trained on clean images, evaluating on denoised predictions better reflects its behavior. Thus, we first predict $\hat{\boldsymbol{x}}_{0}$ from $\bar{\boldsymbol{x}}_{t-1}$, then measure $f(\boldsymbol{x}_{tar}|\hat{\boldsymbol{x}}_{0})$. Derivation is in Supplementary Material 6.2
3.3 ScoreCAM-based Inpainting
We introduce ScoreCAM to guide the diffusion denoising process with a reference image, generating natural adversarial examples that resemble it. Unlike gradient-based methods such as GradCAM [42], which suffer from vanishing gradients and instability [49], ScoreCAM avoids gradient dependence.
Figure 3: Illustration of ScoreCAM.
Thus, for a given class $c$, ScoreCAM can be computed as:
We leverage the concept of image inpainting [31] to perform a weighted fusion of $\boldsymbol{x}_{t-1}^{ref}$ with $\boldsymbol{\tilde{x}}_{t}$.
3.4 Noise Optimization
[67] show that certain noise types yield better results. Motivated by this, we incorporate prior information about the target label into the initial noise. Specifically, after generating $\boldsymbol{x}_{0}$, we apply diffusion process and use gradient signals from the target model to guide noise optimization.
$\boldsymbol{x}_{T}=\left(\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}\right)+\bar{\sigma}_{T}^{2}s_{n}\nabla_{\boldsymbol{x}_{0}}\log p_{f}(y_{tar}|\boldsymbol{x}_{0}),$ (15)
$\boldsymbol{x}_{T}=\left(\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}\right)+\bar{\sigma}_{T}^{2}s_{n}\nabla_{\boldsymbol{x}_{0}}\log p_{f}(\boldsymbol{x}_{tar}|\boldsymbol{x}_{0}).$ (16)
Full derivation is in Supplementary Material 6.3.
3.5 ScoreAdv Algorithm
4 Experiments
4.1 Datasets